Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 203     Entry time: 12 May 2023, 23:21
Author: Romain Letrun 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, 1.8 mJ, 202 pulses most of the day.
  • Attempted tuning to 14.8 keV during the first hour fo the shift did not work out. More time will be spent tomorrow morning tuning to that photon energy.
  • Overall stable delivery. Only a magnet failure at 20:25 resulted in a drop of energy and was quickly recovered. Interestingly, during that time the slope through the train was reduced.

Optical laser delivery

  • NA

Achievements/Observations

  • Collected diffraction data from multiple samples, including mixing at different delays. Smooth sailing with the same nozzle for the entire day.

Issues

  • Run controller and data manager crashed around 23:00 just as we were about to collect data from the last sample. Both devices were reinstantiated.
ELOG V3.1.4-7c3fd00